# Floods do not sink prices, historical memory does: How flood risk impacts the Italian housing market

### v1

@@ 2.1 Intesa Sanpaolo mortgages
Intesa Sanpaolo (ISP) is the largest banking group in Italy, financing about a quarter of all mortgages within the country. [^11] https://www.agipress.it/mutui-vassena-intesa-sanpaolo-leader-in-italia-con-il-26-del-mercato Following a GDPR-compliant and privacy-preserving protocol, we got access to the vast majority of the mortgages that they issued from 2016 on. Our dataset contains all mortgages that are still being repaid at the time of data extraction, that is August 2024.121212This excludes all the mortgages that were issued but have already been fully repaid, and mortgages that were issued but have been subrogated. However, focusing on mortgages that are less than 8 years old, these cases are not very common and we estimate that even in the initial year of the sample (2016-2018) we have 80% of the mortgages that were actually issued. The raw dataset contains 798,237 rows. We filtered data by selecting mortgages used to finance purchases (rather than, for instance, renovations), with available coordinates and within the Italian territory, and following a few other criteria. [^13] We also discarded the mortgages that were given to juridical persons, that are still under review, that were issued to construct housing units with the goal of reselling them, and that were acquired through an auction process. The final dataset has 552,761 rows.
@@ 2.3 Copernicus flooded areas
2.4 Historical memory: frequency of floods across Italy
Our hypothesis is that historical memory is built through the repeated exposure to floods, not after single events. Therefore, to quantify the level of historical memory of the different Italian regions we relied on historic flood records reported in the EM-DAT (2024) database.
@@ 2.4 Historical memory: frequency of floods across Italy
EM-DAT is a disaster database that contains all the floods that have occurred in Italy with at least one among: 10 fatalities; 100 affected people; a declaration of state of emergency; a call for international assistance. These objective inclusion criteria and the popularity of the database make it an ideal starting point for our analysis.
We decided to consider the simple count of flood events by region as an indicator of how frequently a region is affected by floods, disregarding their severity. [^20] EM-DAT reports multiple severity measures for each event, e.g. fatalities, missing and displaced people, total damage, insured damage, etc. However, it aggregates this information by events that typically span multiple regions, making it difficult to apportion damages to each region. Moreover, for many floods some of this information is missing. This allowed us to identify 63 event-region pairs in the 2000-2023 period [^21] EM-DAT discourages using data before 2000 because of reliability concerns., Figure 1 shows them disaggregated by region.
These flood records allowed us to classify the exposure of Italian regions into 3 different levels of historical memory: Frequently flooded regions, Moderately flooded regions, and Rarely flooded regions, which correspond to the 3 terciles of the distribution shown in Figure 1. A map of the regions associated to each historical memory level can be found in Figure 4.
We also considered as a robustness two alternative measures of flood frequency: (1) we normalize the number of events on EM-DAT by region area, (2) we consider an alternative dataset that we compiled from ISPRA data. Supplementary Section B.1 provides more details.
Figure 1: Flood events identified by EM-DAT, disaggregated by region, and colored by macro-area. 8/12 events in Veneto are likely to correspond to the acqua alta (high-tide) events that affect the city of Venice, as only Venice is mentioned by EM-DAT in these events. Without considering these events, Emilia-Romagna and Toscana are the most frequently flooded regions.
@@ 3 Econometric design
To quantify the influence of flood risk on the Italian housing market, we have addressed three separate research questions: (1) Do we observe changes in home prices after specific flood events?; (2) Does flood risk affect home prices differently in different regions?; (3) How do results change across socio-economic groups? In the following, we describe how we address all these questions by building on a baseline hedonic regression.
@@ 3.1 Baseline hedonic regression
where $P_{ikt}$ is the transaction price, $\alpha_{k}$ denotes spatial fixed effects, $\alpha_{t}$ denotes time fixed effects, $R_{i}$ is a dummy variable that indicates that a home is at flood risk according to ISPRA maps, and $\textbf{X}_{i}$ is a vector of property characteristics. We now discuss each of these terms one by one.
In our formulation, $\alpha_{k}$ captures all time-invariant home price differences due to being located in spatial unit $k$. We consider different definitions of spatial units, at different spatial aggregation levels. The widest definition of $k$ is the municipality, with a median area of 22 km2. Especially for large municipalities, it is unlikely that this fixed effect captures enough difference in location, so we consider two further narrower definitions of $\alpha_{k}$. These are the OMI microzone and the census tract, respectively featuring a median area of 1.5 km2 and 0.03km2.222222The OMI microzone is the standard spatial unit of the Osservatorio del Mercato Immobiliare (OMI), the housing market observatory of the Italian Tax Office. It roughly corresponds to a neighborhood, although in rural areas it may contain larger portions of territory. Instead, census tracts correspond to blocks in cities and to villages (“frazioni”) in rural areas. We use the latest available census tracts, namely from the 2011 census. The inclusion of the $\alpha_{k}$ term completely captures the spatial correlation existing in our data, see Supplementary Section B.2. This confirms that our hedonic regression approach is adequate and spatial econometric approaches are not necessary.
@@ 3.2 Effect of flood events
$\quad\delta_{1}\cdot\texttt{T}_{i}\times\texttt{HitRisk}_{i}+\delta_{2}\cdot\texttt{T}_{i}\times\texttt{NoHitRisk}_{i}+\boldsymbol{\gamma}^{\prime}\textbf{X}_{i}+\epsilon_{ikt},$ (2)
The categorical variable T reports the binned time of the transaction compared to the flood, using the period before the flood as a reference. In this specification, the coefficients that we are mostly interested in are the ones on the interaction terms, namely $\delta_{1}$ and $\delta_{2}$, capturing, respectively, the additional reduction in prices for homes at risk in the flood affected area, and the additional reduction for homes not directly hit by the flood but still considered at risk. In short, $\delta_{1}$ captures the direct effect due to being hit, whereas $\delta_{2}$ captures the indirect effect due to the beliefs of the market.
@@
3.3 Spatially disaggregated impact of flood risk
@@
First, we dis-aggregate the effects of buying a home at risk by frequency of past flood events, i.e., what we call historical memory. To this aim, we interact the risk indicator with a categorical indicator HistMem detailing the home region’s historical level of exposure to floods: frequently flooded, moderately flooded or rarely flooded (see Section 2.4 for a more detailed definition). This results in the following equation
$\log{P_{ikt}}=\alpha_{k}+\alpha_{t}+\rho R_{i}\times\texttt{HistMem}_{i}+\boldsymbol{\gamma}^{\prime}\textbf{X}_{i}+\epsilon_{ikt}.$ (3)
Next, at a more granular level, we also distinguish across specific regions, replacing the HistMem categorical variable by a region variable.
@@ 3.4 Disaggregation by socio-economic groups
We encoded the age of the buyer in the variable Age with two categories: Young and NotYoung. This is based on whether the mortgage was a mutuo giovani (Young Buyers’ Mortgage) or not. Transactions financed by the mutuo giovani are associated with buyers under 36 with an adjusted annual income lower than 40,000€. We assumed all mortgages not labeled in the data as mutuo giovani to refer to NotYoung buyers.
For the level of income of the buyers, we divided buyers into 3 levels of income (high, medium, and low) based on the three terciles of the income distribution for each region. We define income levels by region because average income is widely different across Northern and Southern regions, and we do want this variable to capture relative income within regions.
@@
By performing the triple interaction between the $R_{i}$ indicator, the HistMem and Age variables, we can estimate whether young or more experienced buyers living in regions with different levels of historical flood exposure obtain different price reductions when buying homes at risk. The equation used for this analysis is the following:
$\log{P_{ikt}}=\alpha_{k}+\alpha_{t}+\tau R_{i}\times\texttt{HistMem}_{i}\times\texttt{Age}_{i}+I+\boldsymbol{\gamma}^{\prime}\textbf{X}_{i}+\epsilon_{ikt}.$ (4)
In the above equation, the key coefficient is $\tau$, capturing the effect of risk on home prices for each combination of age and historical flood frequency. In this regression we also included the term $I$ to control for buyers’ income, accounting for income differences between Young and NotYoung buyers (12%).
@@
$\log{P_{ikt}}=\alpha_{k}+\alpha_{t}+\psi R_{i}\times\texttt{HistMem}_{i}\times\texttt{Age}_{i}\times\texttt{IncomeLevel}_{i}+I+\boldsymbol{\gamma}^{\prime}\textbf{X}_{i}+\epsilon_{ikt}$ (5)
where the coefficient $\psi$ indicates the disaggregated effect of age, level of income, and historical memory, on the price of homes at risk. Thus, it makes it possible to distinguish the reduction obtained by, e.g., low-income and high-income buyers that are both young.
@@ 4.1 Specific floods only affect home prices in flooded areas
Figure 3: Time evolution of the home price variation for the 2023 Emilia-Romagna flood. Referring to Equation (2), the left panel reports the coefficients $\beta_{1}$ and $\beta_{2}$, the central and right panels show the interaction coefficients $\delta_{1}$ and $\delta_{2}$ up to 15 months after the flood of Emilia-Romagna 2023. The temporal bins can be interpreted as follows. The labels pre and post identify whether a transaction has been completed before or after the considered flood event. The label y indicates years and the label m indicates months. We used the bin pre 1y as a reference. Each point represents the average value of the regression coefficient and the error bars indicate the 95% confidence intervals. R2 = 0.74, observations = 35623.
@@
Figure 3 shows the results. The left panel shows the baseline discount for being in a risk zone for homes that were hit ($\beta_{1}$) and that were not hit ($\beta_{2}$). These coefficients state that HitRisk and NoHitRisk homes sell –on average– around 3% less than homes in safe areas. Further details about this results will be provided in the next subsections.
Next, focusing on homes that were hit (central panel), we find no trend prior to the event, virtually no price reduction in the first 3 months after the event, and then a substantial reduction between 3 and 9 months from the event. In this period, homes sell at a further 10% less than the baseline 3% for homes at risk. However, in the last two quarters of observations, the prices actually increase compared to the baseline: homes that were hit by the flood sell at a 10% premium compared to homes that were not hit (and not in the risk zone). We interpret these results as follows. In the first three months after the event, prices did not decrease because many transactions reflect deals that were agreed upon before the flood (with possibly a pre-authorization on the mortgage amount by the bank). In the following months, the reductions in prices reflect the cost of the renovation work that falls on the buyer. Finally, 9 months after the event most renovation work has been carried out, and so the price increase reflects newly renovated homes [^26] We unfortunately do not have data on renovation status to test this hypothesis, but it could be tested in the future as ISP is increasingly collecting this information. We think this is a reasonable timeline for renovation for the type of damage suffered by homes in this flood.
@@
These results were also consistent when considering alternative parameter specifications, when analyzing only ground-floor apartments or independent houses, and when exploring other major flood events (see Supplementary Section B.1).
4.2 Homes at risk sell at lower price in frequently flooded regions (“historical memory”)
In the left panel of Figure 3, we showed that being located in an at-risk area leads to a lower home price, irrespective of the specific flood considered. To explore this question further, we estimate Equation (1) on our dataset aiming to provide a national-scale picture of flood risk across Italy.
Table 2: Flood risk effect on home prices. For each spatial fixed effects, Models 1-3-5 report the effect of the risk on home prices estimated using Equation (1), while Models 2-4-6 dis-aggregate the effect of flood risk by the historical memory of the regions according to EM-DAT (2024), Equation (3). *** $p<0.01$, ** $p<0.05$, * $p<0.1$.

### v2

@@ 2.1 Intesa Sanpaolo mortgages
Intesa Sanpaolo (ISP) is the largest banking group in Italy, financing about a quarter of all mortgages within the country. [^11] https://www.agipress.it/mutui-vassena-intesa-sanpaolo-leader-in-italia-con-il-26-del-mercato Following a GDPR-compliant and privacy-preserving protocol, we got access to the vast majority of the mortgages that they issued from 2016 on. Our dataset contains all mortgages that are still being repaid at the time of data extraction, that is August 2024.121212This excludes all the mortgages that were issued but have already been fully repaid, and mortgages that were issued but have been subrogated. However, focusing on mortgages that are less than 8 years old, these cases are not very common and we estimate that even in the initial year of the sample (2016-2018) we have 80% of the mortgages that were actually issued. The raw dataset contains 798,237 rows. We filtered data by selecting mortgages used to finance purchases (rather than, for instance, renovations), with available coordinates and within the Italian territory [^13] Homes without available coordinates had to be removed from the analysis due to the impossibility of classifying them into risk areas., and following a few other criteria. [^14] We also discarded the mortgages that were given to juridical persons, that are still under review, that were issued to construct housing units with the goal of reselling them, and that were acquired through an auction process. The final dataset has 552,761 rows.
@@ 2.3 Copernicus flooded areas
2.4 Flood risk awareness: frequency of floods across Italy
To model buyers’ awareness of flood risk we combined historical data of flood events in Italian regions with the existing theoretical insights on flood awareness perception.
@@ 2.4 Flood risk awareness: frequency of floods across Italy
Historical data on flood events come from EM-DAT, which is a disaster database that contains all the floods that have occurred in Italy between 2000 and 2023212121Though EM-DAT provides older event records, it discourages the use of the pre-2000 data because of reliability concerns. with at least one among: 10 fatalities; 100 affected people; a declaration of state of emergency; a call for international assistance. These objective inclusion criteria and the popularity of the database make it an ideal starting point for our analysis. As a robustness test, we also considered an alternative dataset that we compiled from ISPRA data (Supplementary Section B.1 provides more details). In both cases, geographical information on floods is provided at the region level, making regions our natural spatial unit for computing awareness over time.
From the theoretical literature on flood risk awareness it is known that perception of risk tends to rise after a new catastrophic event but then it slowly decays as the event is forgotten (Pryce et al., 2011; Tobin and Newton, 1986). Some studies also measure the time scale of awareness. For instance, Heimerl and Meyer (2014) and Egli (2002) report that awareness halves after 7 to 10 years, while Barendrecht et al. (2019) find a 17 years half-life (i.e., the time it takes for a quantity to halve).
We integrate these insights and the actual records of past flood events from EM-DAT to create our estimate of the flood awareness for each Italian region (we cannot go to a lower level of granularity because information on provinces is often lacking or inconsistent). We define the awareness to flood risk in region $r$ at date $t$ as
$a(r,t)=\sum_{d=01/01/2000}^{t}e^{-\frac{(t-d)\cdot\ln{2}}{365\cdot\tau}}\cdot f(r,d),$ (1)
where $f(r,d)$ is 1 if there was a flood in region $r$ at day $d$ according to EM-DAT and 0 otherwise, [^22] If the flood event spans multiple days, we consider the initial day as $d$. We did not consider event severity in our analysis. Although EM-DAT reports multiple severity measures for most events (e.g. fatalities, missing and displaced people, total damage, insured damage, etc), it aggregates this information by events that typically span multiple regions, and this makes it difficult to apportion damages to each region. and $\tau$ is the half-life parameter that represents the time required to halve the influence of each flood event in $a(r,t)$. Thus, the awareness of an Italian region at time $t$ is a weighted sum of the events it experienced, but the contribution of each event decays exponentially as the gap between $t$ and the event time $d$ gets larger. In this way, recent floods contribute more, but their influence decays over time as the event is forgotten.
Overall, $a(r,t)$ combines in a unique variable the number of floods a region has experienced (spatial information) with the time passed between the past floods and the transaction date (temporal information) giving us a good estimate of buyers’ flood risk awareness.
Figure 1 shows the estimated awareness trends per region for the period covered by the ISP data, January 1st, 2016 to August 2024. As the reader can see, starting from 2016, regions show different awareness levels and trends due to the different exposure to floods between 2000 and 2016. For example, we can compare the Emilia-Romagna and Puglia regions. Emilia-Romagna shows a 2.5 awareness level in 2016 and experienced several floods up to August 2024 (spikes in 2018, middle 2019, 2021, and middle 2023) that explain the positive awareness trend. In contrast, Puglia was hit by only one flood in 2003, leading to a low awareness value that decreases in time due to the exponential decay of the weight (Equation 1). [^23] In the main text, we show results considering $\tau=10$ years. 10 years is a value in line with a previous study on the perception of flood risk (Barendrecht et al., 2019). Table A7 in reports the results considering $\tau=7$ and 17 years (coherently with (Barendrecht et al., 2019)), highlighting that our results are robust across different value of $\tau$. A map showing the average awareness of the transaction within each region can be found in Figure 4.
Figure 1: Regional awareness trends computed on the EM-DAT data. The awareness is computed day-by-day according to Equation (1) with $\tau=10$ years. The spikes are due to flood events.
Building on Equation (1), we associate with each transaction in the ISP dataset its awareness at the time of the sale $t_{s}$. In other words, $a(r_{s},t_{s})$ quantifies the level of exposure to past floods that the buyer of each home had experienced when they decided to buy it.
@@ 3 Econometric design
To quantify the influence of flood risk on the Italian housing market, we have addressed three separate research questions: (1) Do we observe changes in home prices after specific flood events?; (2) Does flood risk affect home prices? Is this influenced by repeated exposure to floods (historical memory)?; (3) How do results change across socio-economic groups? In the following, we describe how we address all these questions by building on a baseline hedonic regression.
@@ 3.1 Baseline hedonic regression
where $P_{ikt}$ is the transaction price, $\alpha_{k}$ denotes spatial fixed effects, $\alpha_{t}$ denotes time fixed effects, $R_{i}$ is a dummy variable that indicates that a home is at flood risk according to ISPRA maps, [^24] Although it is unclear to what extent ISPRA flood maps are known to prospective buyers (given that they are publicly available but, as discussed in the introduction, flood risk information is not formally disclosed in Italy) we assume that these maps reflect an intuitive notion of flood risk. For example, proximity to a river is an obvious indicator of flood exposure, and such features are indeed captured in the ISPRA maps. Therefore, we use ISPRA classifications as a proxy for areas that buyers are likely to perceive as at risk. and $\textbf{X}_{i}$ is a vector of property characteristics. We now discuss each of these terms one by one.
In our formulation, $\alpha_{k}$ captures all time-invariant home price differences due to being located in spatial unit $k$. We consider different definitions of spatial units, at different spatial aggregation levels. The widest definition of $k$ is the municipality, with a median area of 22 km2. Especially for large municipalities, it is unlikely that this fixed effect captures enough difference in location, so we consider two further narrower definitions of $\alpha_{k}$. These are the OMI microzone and the census tract, respectively featuring a median area of 1.5 km2 and 0.03km2.252525The OMI microzone is the standard spatial unit of the Osservatorio del Mercato Immobiliare (OMI), the housing market observatory of the Italian Tax Office. It roughly corresponds to a neighborhood, although in rural areas it may contain larger portions of territory. Instead, census tracts correspond to blocks in cities and to villages (“frazioni”) in rural areas. We use the latest available census tracts, namely from the 2011 census. The inclusion of the $\alpha_{k}$ term completely captures the spatial correlation existing in our data, see Supplementary Section B.2. This confirms that our hedonic regression approach is adequate and spatial econometric approaches are not necessary.
@@ 3.2 Effect of flood events
$\quad\delta_{1}\cdot\texttt{T}_{i}\times\texttt{HitRisk}_{i}+\delta_{2}\cdot\texttt{T}_{i}\times\texttt{NoHitRisk}_{i}+\boldsymbol{\gamma}^{\prime}\textbf{X}_{i}+\epsilon_{ikt}.$ (3)
The categorical variable T reports the binned time of the transaction compared to the flood, using the period before the flood as a reference. In this specification, the coefficients that we are mostly interested in are the ones on the interaction terms, namely $\delta_{1}$ and $\delta_{2}$, capturing, respectively, the additional reduction in prices for homes at risk in the flood affected area, and the additional reduction in prices for homes not directly hit by the flood but still considered at risk. In short, $\delta_{1}$ captures the direct effect due to being hit, whereas $\delta_{2}$ captures the indirect effect due to the beliefs of the market.
@@
3.3 Disaggregated impacts of flood risk
@@
First, we disaggregate the effects of buying a home at risk by buyers’ flood awareness at the time of the transaction, i.e. our measure of historical memory derived from historical flood records (Section 2.4). To this aim, we interact the risk indicator with a categorical indicator Awareness detailing the level of awareness in the region at the time of sale. As categories, we consider the terciles of the sale-time awareness distribution across all Italy, and name them high, medium and low awareness. This results in the following equation
$\log{P_{ikt}}=\alpha_{k}+\alpha_{t}+\rho R_{i}\times\texttt{Awareness}_{i}+\boldsymbol{\gamma}^{\prime}\textbf{X}_{i}+\epsilon_{ikt}.$ (4)
Next, we also disaggregate the effects of buying a home at risk across specific regions, providing a more granular overview of the heterogeneous impacts of flood risk across Italy.
@@ 3.4 Disaggregation by socio-economic groups
We encoded the age of the buyer in the variable Age with two categories: Young and NotYoung. This is based on whether the mortgage was a mutuo giovani (Young Buyers’ Mortgage) or not. Transactions financed by the mutuo giovani are associated with buyers under 36 with an adjusted annual income lower than 40,000€. We assumed all mortgages not labeled in the data as mutuo giovani to refer to NotYoung buyers. [^28] Note that only individuals under the age of 36 are eligible to apply for a mutuo giovani. Conversely, while it is possible for a young buyer to obtain a standard mortgage, this is uncommon due to the more favorable terms offered by the mutuo giovani.
For the level of income of the buyers, we divided buyers into 3 levels of income (high, medium, and low) based on the three terciles of the income distribution for each region. We define income levels by region because average income is widely different across Northern and Southern regions, and we want this variable to capture relative income within regions.
@@
By performing the triple interaction between the $R_{i}$ indicator, the Awareness and Age variables, we can estimate whether young or more experienced buyers obtain different price reductions when buying homes at risk under different flood awareness levels. The equation used for this analysis is the following:
$\log{P_{ikt}}=\alpha_{k}+\alpha_{t}+\zeta R_{i}\times\texttt{Awareness}_{i}\times\texttt{Age}_{i}+I+\boldsymbol{\gamma}^{\prime}\textbf{X}_{i}+\epsilon_{ikt}.$ (5)
In the above equation, the key coefficient is $\zeta$, capturing the effect of risk on home prices for each combination of age and awareness. In this regression we also included the term $I$ to control for buyers’ income, accounting for income differences between Young and NotYoung buyers (12%).
@@
$\log{P_{ikt}}=\alpha_{k}+\alpha_{t}+\psi R_{i}\times\texttt{Awareness}_{i}\times\texttt{Age}_{i}\times\texttt{IncomeLevel}_{i}+I+\boldsymbol{\gamma}^{\prime}\textbf{X}_{i}+\epsilon_{ikt}$ (6)
where the coefficient $\psi$ indicates the disaggregated effect of age, level of income, and flood awareness, on the price of homes at risk. Thus, it makes it possible to distinguish the reduction obtained by, e.g., low-income and high-income buyers that are both young.
@@ 4.1 Specific floods only affect home prices in flooded areas
Figure 3: Time evolution of the home price variation for the 2023 Emilia-Romagna flood. Referring to Equation (3), the left panel reports the coefficients $\beta_{1}$ and $\beta_{2}$, the central and right panels show the interaction coefficients $\delta_{1}$ and $\delta_{2}$ up to 15 months after the flood of Emilia-Romagna 2023. The temporal bins can be interpreted as follows. The labels pre and post identify whether a transaction has been completed before or after the considered flood event. The label y indicates years and the label m indicates months. We used the bin pre 1y as a reference. Each point represents the average value of the regression coefficient and the error bars indicate the 95% confidence intervals. We used spatial fixed effects at the OMI microzone level. R2 = 0.74, observations = 35623.
@@
Figure 3 shows the results. The left panel shows the baseline discount for being in a risk zone for homes that were hit ($\beta_{1}$) and that were not hit ($\beta_{2}$). These coefficients state that HitRisk and NoHitRisk homes sell –on average– around 3% less than homes in safe areas, in line with results reported in the next subsection.
Next, focusing on homes that were hit (central panel), we find no trend prior to the event, virtually no price reduction in the first 3 months after the event, and then a substantial reduction between 3 and 9 months from the event. In this period, homes sell at a further 10% less than the baseline 3% for homes at risk. However, in the last two quarters of observations, the prices actually increase compared to the baseline: homes that were hit by the flood sell at a 10% premium compared to homes that were not hit (and not in the risk zone). [^30] All the quarters post-flood have similar number of transactions, with an average of 363 transactions per quarter. We interpret these results as follows. In the first three months after the event, prices did not decrease because many transactions reflect deals that were agreed upon before the flood (with possibly a pre-authorization on the mortgage amount by the bank). In the following months, the reductions in prices reflect the cost of the renovation work that falls on the buyer. Finally, 9 months after the event most renovation work has been carried out, and so the price increase reflects newly renovated homes. [^31] We unfortunately do not have data on renovation status to test this hypothesis, but it could be tested in the future as ISP is increasingly collecting this information. We think this is a reasonable timeline for renovation for the type of damage suffered by homes in this flood.
@@
These results were also consistent when considering alternative parameter specifications (such as more granular fixed effects), when analyzing only ground-floor apartments or independent houses, and when exploring other major flood events (see Supplementary Section B.1). [^33] Our study on the 2023 Emilia-Romagna flood (EMS154) can be consistently compared to previous studies focusing on homes in the 100-year floodplain (Atreya et al., 2013; Bin and Landry, 2013). In fact, almost all at-risk homes in the municipalities affected by the Emilia-Romagna flood are in high or medium risk areas; thus, they are characterized by return times less or equal than 200 years. We only find a negligible number of homes placed in the low-risk zones (9 homes in total). To the best of our knowledge, considering ISPRA data, high or medium risk areas are the those that better align with the 100-years flood zones used in previous studies.
4.2 Homes at risk sell at lower price under high awareness conditions
In the left panel of Figure 3, we showed that being located in an at-risk area leads to a lower home price, irrespective of the specific flood considered. To explore this question further, we estimate Equation (2) on our entire dataset aiming to provide a national-scale picture of flood risk across Italy.
Table 2: Flood risk effect on home prices. For each spatial fixed effects, Models 1-3-5 report the effect of the risk on home prices estimated using Equation (2), while Models 2-4-6 disaggregate the effect of flood risk by the level of flood awareness at the time of the purchase, Equation (4). Awareness is estimated following the procedure described in Section 2.4 using the data reported by EM-DAT (2024) and a half life time equal to 10 years. *** $p<0.01$, ** $p<0.05$, * $p<0.1$.